# Uncertainty, Single Crossing Property, and Stochastic Choice Data

### v1

@@ 3 Proof of Necessity
For the rest of this section, let $u$ be a single crossing utility function, $\mathcal{I}$ be an MLR-ordered information structure, $\pi\in\Delta(\Delta(\Theta))$ be the distribution over posteriors induced by it, and $C:\text{supp}(\pi)\rightarrow\Delta(A)$ be a choice rule. Thus, the tuple $\langle u,\mathcal{I},C\rangle$ corresponds to a Bayesian decision maker.
Since $\mathcal{I}$ is MLR-ordered and $u$ is single crossing, MCS must hold. For any $\gamma\in\text{supp}(\pi)$, define the support of $C(\gamma)$ as $\text{supp}(C(\gamma)):=\{a\in A|C(a|\gamma)>0\}$. Rationality requires that each action in the support of the choice rule must be optimal. Therefore, the support must exactly be solution set to the DM’s maximization problem. MCS then requires that the solution must be ordered by the strong set order. The following establishes this argument formally.
@@
Suppose, for now, that the choice rule is degenerate at each belief. That is, $C(\gamma)=\delta_{a}$ for some $a$. By Lemma 1, $C$ is increasing in the sense that if $\gamma^{\prime}\succ_{MLR}\gamma$, $C(\gamma^{\prime})=\delta_{a^{\prime}}$, and $C(\gamma)=\delta_{a}$, then $a^{\prime}\geq a$ [^13] Equivalently, in this case, $C(\gamma^{\prime})$ first order stochastically dominates $C(\gamma)$. For each $a\in A$, define $\Gamma(a):=\{\gamma\in\text{supp}(\pi)|C(\gamma)=\delta_{a}\}$. That is, $\Gamma(a)$ is the set of all posteriors at which $a$ is optimal. Then, the probability of observing action $a$ in state $\theta$ is given by $\tilde{q}(a|\theta):=\sum_{\gamma\in\text{supp}(\pi)}\pi(\gamma|\theta)C(a|\gamma)=\sum_{\gamma\in\Gamma(a)}\pi(\gamma|\theta)$. Obviously, data matching requires that $\tilde{q}(a|\theta)$ coincides with the observed data. Finally, let $\hat{A}:=\{a\in A|\exists\gamma\in\text{supp}(\pi)\text{ such that }C(\gamma)=\delta_{a}\}$ so that $\hat{A}$ is the set of actions which are optimal for some posterior. Put another way, $\hat{A}$ is the set of actions that may be (optimally) taken by the DM.
@@
For the case where the choice rule is deterministic - so that for each posterior there is a unique optimal action - Lemma 2 establishes this argument formally. In this case, the result rests upon the observation that the posteriors which induce a higher action must be higher - in the the sense of MLR - than those that induce a lower action. This follows from the contrapositive of MCS and our assumption that the sets of signals, states, and actions are strictly linearly ordered. Suppose $\theta_{2}>\theta_{1}$ and $a_{2}>a_{1}$. Then, the MLR property of $\mathcal{I}$ implies that, under $\theta_{2}$, the total probability of obtaining a posterior that induces $a_{2}$ is higher than that of observing a posterior that induces $a_{1}$. The opposite would hold under $\theta_{1}$. Moreover, this holds for all such pairs of states and actions, which essentially means that the collection of conditional distributions over actions is MLR-ordered.
@@
where $s_{\gamma}$ denotes the signal realization that induces the signal $\gamma$ [^14] Note that this is without loss of generality. If, in a given information structure, multiple signal realizations induce same posterior, then they can simply be “collapsed” together. By MCS, $a<a^{\prime}$ implies $s_{\gamma}\leq s_{\gamma^{\prime}}$ for any $\gamma^{\prime}\in\Gamma(a^{\prime})$ and $\gamma\in\Gamma(a)$. As $C(\gamma)$ is degenerate for each $\gamma$, $\Gamma(a)\cap\Gamma(a^{\prime})=\emptyset$. Moreover, as argued above, it is without loss of generality to assume that each signal realization induces a unique posterior, and therefore it must be that $s_{\gamma}<s_{\gamma^{\prime}}$. The latter follows from the assumption that the set of signals is completely ordered by a strict linear order. Now, for $\theta^{\prime}>\theta$, for $\{\tilde{q}(\cdot|\theta)\}_{\theta\in\Theta}$ to be MLR-ordered one needs
$\tilde{q}(a^{\prime}|\theta^{\prime})\tilde{q}(a|\theta)$ $>\tilde{q}(a^{\prime}|\theta)\tilde{q}(a|\theta^{\prime})$
$\Leftrightarrow\left(\sum_{\gamma\in\Gamma(a^{\prime})}\mu(s_{\gamma}|\theta^{\prime})\right)\left(\sum_{\gamma\in\Gamma(a)}\mu(s_{\gamma}|\theta)\right)$ $>\left(\sum_{\gamma\in\Gamma(a^{\prime})}\mu(s_{\gamma}|\theta)\right)\left(\sum_{\gamma\in\Gamma(a)}\mu(s_{\gamma}|\theta^{\prime})\right)$
$\Leftrightarrow\sum_{(\gamma,\gamma^{\prime})\in\Gamma(a)\times\Gamma(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)$ $>\sum_{(\gamma,\gamma^{\prime})\in\Gamma(a)\times\Gamma(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime}).$
@@
Since $\mu(s|\theta)$ is MLR-ordered, any individual term from LHS of the last inequality must be at least as large as any on the RHS. Thus, the only case of concern is one where the number of terms on RHS is greater than those on the LHS. This happens when there are pairs of posteriors $(\gamma,\gamma^{\prime})\in\Gamma(a)\times\Gamma(a^{\prime})$ such that the $\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime})>0$ while $\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)=0$. However, if $\mu(s_{\gamma^{\prime}}|\theta^{\prime})=0$ then $\mu(s_{\gamma^{\prime}}|\theta)$ must be 00 by MLR. Similarly, if $\mu(s_{\gamma}|\theta)=0$ then $\mu(s_{\gamma}|\theta^{\prime})$ must be 00. Therefore, the fact that $\{\mu(\cdot|\theta)\}_{\theta\in\Theta}$ is MLR-ordered ensures that the above inequality holds and thus $\{\tilde{q}(\cdot|\theta)\}_{\theta\in\Theta}$ is MLR-ordered.
@@
Following is the main result of this section. It builds on Lemma 2 and shows that the argument goes through even when the choice rule is non-degenerate - that is, when there are posteriors at which there are multiple optimal solutions. The argument builds on 2 and the fact that the support of choice rule is ordered by the strong set order. Therefore, actions induced by a higher (in the sense of MLR) posterior must dominate - in the strong set order - the actions induced by a lower posterior. Following this observation, the role played by choice rule turns out to redundant, and the MLR inequality is established for the general case.
@@
First note that, if $\gamma^{\prime}\succ_{MLR}\gamma$, it must be that $s_{\gamma^{\prime}}>s_{\gamma}$. For any action $a$, define $\tilde{\Gamma}(a)=\{\gamma\in\text{supp}(\pi)|C(a|\gamma)>0\}$. Consequently, if $a^{\prime}>a$ then $\{s_{\gamma}|\gamma\in\tilde{\Gamma}\}\geq_{sso}\{s_{\gamma}|\gamma\in\tilde{\Gamma}(a)\}_{\theta\in\Theta}$. For $\{q(\cdot|\theta)\}$ to be MLR-ordered, one needs,
@@
$\Leftrightarrow\left(\sum_{\gamma\in\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma}|\theta^{\prime})C(a^{\prime}|\gamma)\right)\left(\sum_{\gamma\in\tilde{\Gamma}(a)}\mu(s_{\gamma}|\theta)C(a|\gamma)\right)$
$>\left(\sum_{\gamma\in\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma}|\theta)C(a^{\prime}|\gamma)\right)$ $\left(\sum_{\gamma\in\tilde{\Gamma}(a)}\mu(s_{\gamma}|\theta^{\prime})C(a|\gamma)\right)$
$\Rightarrow\sum_{(\gamma,\gamma^{\prime})\in\tilde{\Gamma}(a)\times\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)C(a^{\prime}|\gamma^{\prime})C(a|\gamma)$ $>\sum_{(\gamma,\gamma^{\prime})\in\tilde{\Gamma}(a)\times\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime})C(a^{\prime}|\gamma^{\prime})C(a|\gamma)$
In the last inequality, terms corresponding to the case where $\gamma=\gamma^{\prime}$ can be eliminated from both sides. The only remaining terms will thus be the ones where $s_{\gamma^{\prime}}>s_{\gamma}$. Since the terms $C(a^{\prime}|\gamma^{\prime})C(a|\gamma)$ appear on both sides, the result follows from Lemma 2.
@@ 4 Proof of Sufficiency
We now that if the data is MLR ordered, then there exists a single crossing utility function $u$, an MLR ordered information structure $\pi$, and a Choice rule $C$ that rationalize the data. The next subsection shows the construction of the choice rule and information structure.
@@ 4.1 Constructing a single crossing utility function
We proceed to construct a SC utility function inductively. For each consecutive pairs $a+1,a$, we compute single crossing differences $\{A^{\theta}_{a+1,a}\}_{\theta\ in\Theta}$ and then find a utility function with those differences. However, an additional difficulty that needs to be addressed when $N>3$ concerns aggregation. For instance, even if $\{A^{\theta}_{a+1,a}\}$ and $\{A^{\theta}_{a,a-1}\}$ are single crossing, $\{A^{\theta}_{a+1,a-1}\}=\{A^{\theta}_{a+1,a}+A^{\theta}_{a,a-1}\}$ may not be so. That is, the single crossing property is not necessarily preserved under aggregation. However, Quah and Strulovici, 2012 reports the condition - called the signed ratio property - under which the single crossing property is preserved under aggregation. These restrictions turn out to be linear as well, and can therefore be incorporated into the systems of equations that characterize the single crossing differences.
Indeed, these issues arise only when there are more than three states, and the number of restrictions that need to be imposed is thus increasing in the number of states. Before we proceed to inductively construct a single crossing utility function, we note that, for any pair of consecutive actions $a,a+1$, we can pick the state at which the function $\theta\mapsto A^{\theta}_{a+1,a}$ crosses the horizontal axis. In this case, we construct the differences $A^{\theta}_{a+1,a}$ so that the switch happens at $\theta_{n_{a+1}}$.
@@
where $A$ is the first matrix on the left hand side of 2. Suppose that this system does have a solution $\bm{x}=(x_{1},\dots,x_{N+2})\gg 0$. Then, for any $\theta=\theta_{j}<\theta_{n_{a_{k}}},q^{\theta}_{k}x_{1}-q^{\theta}_{k-1}x_{2}-x_{j}=0$. Since $x_{j}>0,q^{\theta}_{k}x_{1}-q^{\theta}_{k-1}x_{2}>0$. That is, $\frac{q^{\theta}_{k}}{q^{\theta}_{k-1}}>\frac{x_{2}}{x_{1}}$. Then, for any $\theta^{\prime}=\theta_{l}\geq\theta_{n_{a_{k}}}>\theta$, $\frac{q^{\theta^{\prime}}_{k}}{q^{\theta^{\prime}}_{k-1}}>\frac{q^{\theta}_{k}}{q^{\theta}_{k-1}}>\frac{x_{2}}{x_{1}}$, where the first inequality follows from the MLR property. Therefore, $q^{\theta^{\prime}}_{k}x_{1}-q^{\theta^{\prime}}_{k-1}x_{2}>0$. But $\bm{x}$ is a solution to the dual system, and so $q^{\theta^{\prime}}_{k}x_{1}-q^{\theta^{\prime}}_{k-1}x_{2}+x_{l}=0$ which is a contradiction since $x_{l}>0$.
Thus, the system $A^{\prime}x=0$ does not have a solution $x\gg 0$. By theorem 2, the system 2 must have a solution. ∎
So, there exist solutions $\{A^{\theta}_{2,1}\}_{\theta=1}^{N}$. Note that because of the last $N-2$ constraints in 2, these two solutions satisfy single crossing by construction.
@@
    For each $l=1,\dots,k-2$, define $\widetilde{A}^{k}=\{a_{l}\in A||\widetilde{\Theta}^{k,l}|\geq 2\}$.
@@
    Let $\bm{\kappa(\widetilde{\Theta}^{k,l})}$ be the matrix obtained by stacking all the row vectors 3 on top of each other.
@@
For $l=k-1,k-2$, SCP follows by construction. For $l<k-2$, and by Claim 2, we only need to take care of cases with $A^{\theta}_{k,k-1}<0,A^{\theta}_{k-1,l}>0$. These are exactly the states $\theta$ that lie in $\widetilde{\Theta}^{k,l}$ and the restrictions involving $\bm{\kappa_{\widetilde{A}_{k}}}$ ensure that the appropriate version of signed ratio property holds for all such action-state pairs. SCP then follows by Proposition 1 of Quah and Strulovici, 2012. ∎
@@
For $l=k-1$, optimality follows by construction. Now consider any $l<k-1$. Recall that $A^{\theta}_{k,l}=A^{\theta}_{k,k-1}+A^{\theta}_{k-1,k-2}+\dots+A^{\theta}_{l+1,l}$. We use induction on $l$ to prove the result. First, consider $A^{\theta}_{k,k-2}=A^{\theta}_{k,k-1}+A^{\theta}_{k-1,k-2}$. By construction, $\sum_{\theta}q^{\theta}_{k}A^{\theta}_{k,k-1}\geq 0,\sum_{\theta}q^{\theta}_{k-1}A^{\theta}_{k,k-1}\leq 0,\sum_{\theta}q^{\theta}_{k-1}A^{\theta}_{k-1,k-2}\geq 0$ and, $\sum_{\theta}q^{\theta}_{k-2}A^{\theta}_{k-1,k-2}\leq 0$. Now consider,
@@
Thus, in 6, replacing the coefficients $\left(q^{\theta}_{k-1}\frac{q^{k-2}_{k-2}}{q^{k-2}_{k-1}}\right)$ with $q^{\theta}_{k-2}$ preserves the inequality, since this weakly decreases the coefficients of positive terms and weakly increases the coefficients of negative terms. Therefore,
@@
Then, in 9, replacing the coefficients $\left(q^{\theta}_{k-1}\frac{q^{k-1}_{k}}{q^{k-1}_{k-1}}\right)$ with $q^{\theta}_{k}$ preserves the inequality, since this weakly increases the coefficients of positive terms and weakly decreases the coefficients of negative terms. Therefore,
@@
Now suppose that $\{A^{\theta}_{k,l}\}$, where $l<k-2$, satisfies optimality. To complete the proof, it suffices to show that $\{A^{\theta}_{k,l-1}\}=\{A^{\theta}_{k,l}\}+\{A^{\theta}_{l,l-1}\}$ satisfies optimality. Again, $\{A^{\theta}_{l,l-1}\}$ satisfies optimality by construction. $\{A^{\theta}_{k,l}\}$ satisfies single crossing, so let $\theta_{n_{k,l}}$ be smallest $\theta$ such that $A^{\theta^{\prime}}_{k,l}\geq 0$ for every $\theta\geq\theta^{\prime}$. That is, $\theta_{n_{k,l}}$ is the crossing point for $A^{\theta^{\prime}}_{k,l}$. Consider,
@@
Finally, consider the following $(N-1)\times N^{2}$ system,
@@
There are exactly $(M-1)N$ independent columns in the coefficient matrix on the left hand side. [^18] Each row on the first matrix on the LHS has exactly two 1’s $N^{2}$ 0’s, where the positions of the 1’s ensure that the solution to the system induces the corresponding differences $A^{\theta}_{k,l}$. This is also the dimension of the vector on the RHS. So, the rank of the augmented matrix is same as that of the coefficient matrix and thus there exists a utility function $u(a,\theta)$ which induces the differences $\{A^{\theta}_{k,l}\}$. This completes the construction of a single crossing utility function that rationalizes the data, thereby completing the proof.
@@
4.2 Rationalizing information structure and choice rule

### v3

@@ 4 Proof of Necessity
For the rest of this section, let $u$ be a single crossing utility function, $\mathcal{I}$ be an MLR-ordered information structure, $\pi\in\Delta(\Delta(\Theta))$ be the distribution over posteriors induced by it, and $C:\text{supp}(\pi)\rightarrow\Delta(A)$ be a choice rule. Thus, the tuple $\langle u,\mathcal{I},C\rangle$ corresponds to an MBEU-DM.
Since $\mathcal{I}$ is MLR-ordered and $u$ is single crossing, MCS must hold. For any $\gamma\in\text{supp}(\pi)$, define the support of $C(\gamma)$ as $\text{supp}(C(\gamma)):=\{a\in A|C(a|\gamma)>0\}$. Rationality requires that each action in the support of the choice rule must be optimal. Therefore, under Assumption 1, the support must exactly be the solution set to the DM’s maximization problem. MCS then requires that the solution must be ordered by the strong set order. The following establishes this argument formally.
@@
Suppose, for now, that the choice rule is degenerate at each belief. That is, for each $\gamma\in\text{supp}(\pi)$, $C(\gamma)=\delta_{a}$ for some $a$. By Lemma 1, $C$ is increasing in the sense that if $\gamma^{\prime}\succ_{MLR}\gamma$, $C(\gamma^{\prime})=\delta_{a^{\prime}}$, and $C(\gamma)=\delta_{a}$, then $a^{\prime}\geq a$. For each $a\in A$, define $\Gamma(a):=\{\gamma\in\text{supp}(\pi)|C(\gamma)=\delta_{a}\}$. That is, $\Gamma(a)$ is the set of all posteriors at which $a$ is optimal. Then, the probability of observing action $a$ in state $\theta$ is given by $\tilde{q}(a|\theta):=\sum_{\gamma\in\text{supp}(\pi)}\pi(\gamma|\theta)C(a|\gamma)=\sum_{\gamma\in\Gamma(a)}\pi(\gamma|\theta)$. Obviously, data matching requires that $\tilde{q}(a|\theta)$ coincides with the observed data. Finally, let $\hat{A}:=\{a\in A|\exists\gamma\in\text{supp}(\pi)\text{ such that }C(\gamma)=\delta_{a}\}$ - so that $\hat{A}$ is the set of actions which are optimal for some posterior. Put another way, $\hat{A}$ is the set of actions that may be (optimally) chosen by the DM.
@@
For the case where the choice rule is deterministic - so that for each posterior there is a unique optimal action - Lemma 2 establishes this argument formally. In this case, the result rests upon the observation that the posteriors which induce a higher action must be higher - in the the sense of MLR - than those that induce a lower action. This follows from the contrapositive of MCS and our assumption that the sets of signals, states, and actions are strictly linearly ordered. Suppose $\theta_{2}>\theta_{1}$ and $a_{2}>a_{1}$. Then, the MLR property of $\mathcal{I}$ implies that, under $\theta_{2}$, the likelihood of obtaining a posterior that induces $a_{2}$ relative to one that induces $a_{1}$ is higher than under $\theta_{1}$. Moreover, this holds for all such pairs of states and actions, which essentially means that the collection of conditional distributions over actions is MLR-ordered.
@@
where $s_{\gamma}$ denotes the signal realization that induces the signal $\gamma$. [^15] Note that this is without loss of generality. If, in a given information structure, multiple signal realizations induce same posterior, then they can simply be “collapsed” together. By MCS, $a<a^{\prime}$ implies $s_{\gamma}\leq s_{\gamma^{\prime}}$ for any $\gamma^{\prime}\in\Gamma(a^{\prime})$ and $\gamma\in\Gamma(a)$. As $C(\gamma)$ is degenerate for each $\gamma$, $\Gamma(a)\cap\Gamma(a^{\prime})=\emptyset$. Moreover, as argued above, it is without loss of generality to assume that each signal realization induces a unique posterior, and therefore it must be that $s_{\gamma}<s_{\gamma^{\prime}}$. The latter follows from the assumption that the set of signals is completely ordered by a strict linear order. Now, if $\theta^{\prime}>\theta$, for $\{\tilde{q}(\cdot|\theta)\}_{\theta\in\Theta}$ to be MLR-ordered one needs
$\tilde{q}(a^{\prime}|\theta^{\prime})\tilde{q}(a|\theta)$ $\geq\tilde{q}(a^{\prime}|\theta)\tilde{q}(a|\theta^{\prime})$
$\Leftrightarrow\left(\sum_{\gamma\in\Gamma(a^{\prime})}\mu(s_{\gamma}|\theta^{\prime})\right)\left(\sum_{\gamma\in\Gamma(a)}\mu(s_{\gamma}|\theta)\right)$ $\geq\left(\sum_{\gamma\in\Gamma(a^{\prime})}\mu(s_{\gamma}|\theta)\right)\left(\sum_{\gamma\in\Gamma(a)}\mu(s_{\gamma}|\theta^{\prime})\right)$
$\Leftrightarrow\sum_{(\gamma,\gamma^{\prime})\in\Gamma(a)\times\Gamma(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)$ $\geq\sum_{(\gamma,\gamma^{\prime})\in\Gamma(a)\times\Gamma(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime}).$
@@
Since $\mu(s|\theta)$ is MLR-ordered, any individual term from left side of the last inequality must be at least as large as any on the right. Thus, the only case of concern is one where the number of terms on RHS is greater than those on the LHS. This happens when there are pairs of posteriors $(\gamma,\gamma^{\prime})\in\Gamma(a)\times\Gamma(a^{\prime})$ such that the $\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime})>0$ while $\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)=0$. However, if $\mu(s_{\gamma^{\prime}}|\theta^{\prime})=0$ then $\mu(s_{\gamma^{\prime}}|\theta)$ must be 0 by MLR. Similarly, if $\mu(s_{\gamma}|\theta)=0$ then $\mu(s_{\gamma}|\theta^{\prime})$ must be 0. Therefore, the fact that $\{\mu(\cdot|\theta)\}_{\theta\in\Theta}$ is MLR-ordered ensures that the above inequality holds and thus $\{\tilde{q}(\cdot|\theta)\}_{\theta\in\Theta}$ is MLR-ordered.
@@
Following is the main result of this section. It builds on Lemma 2 and shows that the argument goes through even when the choice rule is non-degenerate - that is, when there are posteriors at which there are multiple optimal solutions. The argument uses 2 and the fact that the support of the choice rule is ordered by the strong set order. Therefore, actions induced by a higher (in the sense of MLR) posterior must dominate - in the strong set order - the actions induced by a lower posterior. Following this observation, the role played by choice rule turns out to be redundant, and the MLR inequality is established for the general case.
@@
First note that, if $\gamma^{\prime}\succ_{MLR}\gamma$, it must be that $s_{\gamma^{\prime}}>s_{\gamma}$. For any action $a$, define $\tilde{\Gamma}(a)=\{\gamma\in\text{supp}(\pi)|C(a|\gamma)>0\}$. Consequently, if $a^{\prime}>a$ then $\{s_{\gamma^{\prime}}|\gamma^{\prime}\in\tilde{\Gamma}(a^{\prime})\}\geq_{sso}\{s_{\gamma}|\gamma\in\tilde{\Gamma}(a)\}_{\theta\in\Theta}$. For $\{q(\cdot|\theta)\}$ to be MLR-ordered, one needs,
@@
$\Leftrightarrow\left(\sum_{\gamma^{\prime}\in\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta^{\prime})C(a^{\prime}|\gamma^{\prime})\right)\left(\sum_{\gamma\in\tilde{\Gamma}(a)}\mu(s_{\gamma}|\theta)C(a|\gamma)\right)$
$>\left(\sum_{\gamma^{\prime}\in\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta)C(a^{\prime}|\gamma^{\prime})\right)$ $\left(\sum_{\gamma\in\tilde{\Gamma}(a)}\mu(s_{\gamma}|\theta^{\prime})C(a|\gamma)\right)$
$\Leftrightarrow\sum_{(\gamma,\gamma^{\prime})\in\tilde{\Gamma}(a)\times\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)C(a^{\prime}|\gamma^{\prime})C(a|\gamma)$ $>\sum_{(\gamma,\gamma^{\prime})\in\tilde{\Gamma}(a)\times\tilde{\Gamma}(a^{\prime})}\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime})C(a^{\prime}|\gamma^{\prime})C(a|\gamma)$
In the last inequality, terms corresponding to the case where $\gamma=\gamma^{\prime}$ can be eliminated from both sides. The only remaining terms will thus be the ones where $s_{\gamma^{\prime}}>s_{\gamma}$. For each of these pairs, the terms $\mu(s_{\gamma^{\prime}}|\theta^{\prime})\mu(s_{\gamma}|\theta)$ are greater than the terms $\mu(s_{\gamma^{\prime}}|\theta)\mu(s_{\gamma}|\theta^{\prime})$ by MLR property and Lemma 2, and the result follows. ∎
@@ 5 Proof of Sufficiency
We now show that if the data is MLR ordered, then there exists a single crossing utility function $u$, an MLR ordered information structure $\pi$, and a choice rule $C$ that rationalize the data. The next subsection shows the construction of the choice rule and information structure.
@@ 5.1 Constructing a single crossing utility function
We proceed to construct a SC utility function inductively. For each consecutive pair $a+1,a$, we compute single crossing differences $\{A^{\theta}_{a+1,a}\}_{\theta\in\Theta}$ and then find a utility function with those differences. However, an additional difficulty that needs to be addressed when $N>3$ concerns aggregation. For instance, even if $\{A^{\theta}_{a+1,a}\}$ and $\{A^{\theta}_{a,a-1}\}$ are single crossing, $\{A^{\theta}_{a+1,a-1}\}=\{A^{\theta}_{a+1,a}+A^{\theta}_{a,a-1}\}$ may not be so. That is, the single crossing property is not necessarily preserved under aggregation. However, Quah and Strulovici, 2012 report the condition - called the signed ratio property - under which the single crossing property is preserved under aggregation. These can be written as linear restrictions and incorporated into the systems of equations that characterize the single crossing differences.
Indeed, these issues arise only when there are more than three states, and the number of restrictions that need to be imposed is increasing in the number of states. Before we proceed to inductively construct a single crossing utility function, we note that, for any pair of consecutive actions $a,a+1$, we can pick the state at which the function $\theta\mapsto A^{\theta}_{a+1,a}$ crosses the horizontal axis. In this case, we construct the differences $A^{\theta}_{a+1,a}$ so that the switch happens at $\theta_{n_{a+1}}$.
@@
where $A$ is the first matrix on the left hand side of 3. Suppose that this system does have a solution $\bm{x}=(x_{1},\dots,x_{N+2})\gg 0$. Then, for any $\theta=\theta_{j}<\theta_{n_{a_{k}}},q^{\theta}_{k}x_{1}-q^{\theta}_{k-1}x_{2}-x_{j}=0$. Since $x_{j}>0,q^{\theta}_{k}x_{1}-q^{\theta}_{k-1}x_{2}>0$. That is, $\frac{q^{\theta}_{k}}{q^{\theta}_{k-1}}>\frac{x_{2}}{x_{1}}$. Then, for any $\theta^{\prime}=\theta_{l}\geq\theta_{n_{a_{k}}}>\theta$, $\frac{q^{\theta^{\prime}}_{k}}{q^{\theta^{\prime}}_{k-1}}>\frac{q^{\theta}_{k}}{q^{\theta}_{k-1}}>\frac{x_{2}}{x_{1}}$, where the first inequality follows from the MLR property. Therefore, $q^{\theta^{\prime}}_{k}x_{1}-q^{\theta^{\prime}}_{k-1}x_{2}>0$. But $\bm{x}$ is a solution to the dual system, and so $q^{\theta^{\prime}}_{k}x_{1}-q^{\theta^{\prime}}_{k-1}x_{2}+x_{l}=0$ - which is a contradiction since $x_{l}>0$.
Thus, the system $A^{\prime}x=0$ does not have a solution $x\gg 0$. By theorem 3, the system 3 must have a solution. ∎
So, there exist solutions $\{A^{\theta}_{2,1}\}_{\theta=1}^{N}$. Note that because of the last N2N-2 constraints in 3, these two solutions satisfy single crossing by construction.
@@
    For each $l=1,\dots,k-2$, define $\widetilde{A}^{k}=\{a_{l}\in A||\widetilde{\Theta}^{k,l}|\geq 2\}$ (these are the only states for the pair $k,l$, for which the aggregation issue is non-trivial).
@@
    (this is essentially the signed ratio property of Quah and Strulovici which preserves SCP under aggregation).
@@
    Let $\bm{\kappa(\widetilde{\Theta}^{k,l})}$ be the matrix obtained by stacking all the row vectors 4 on top of each other.
@@
For l=k1,k2l=k-1,k-2, SCP follows by construction. For l<k2l<k-2, and by Claim 2, we only need to take care of cases with $A^{\theta}_{k,k-1}<0,A^{\theta}_{k-1,l}>0$. These are exactly the states $\theta$ that lie in $\widetilde{\Theta}^{k,l}$ and the restrictions involving $\bm{\kappa_{\widetilde{A}_{k}}}$ ensure that the appropriate version of signed ratio property holds for all such action-state pairs. SCP then follows by Proposition 1 of Quah and Strulovici, 2012. ∎
@@
For l=k1l=k-1, optimality follows by construction. Now consider any l<k1l<k-1. Recall that $A^{\theta}_{k,l}=A^{\theta}_{k,k-1}+A^{\theta}_{k-1,k-2}+\dots+A^{\theta}_{l+1,l}$. We use induction on $l$ to prove the result. First, consider $A^{\theta}_{k,k-2}=A^{\theta}_{k,k-1}+A^{\theta}_{k-1,k-2}$. By construction, $\sum_{\theta}q^{\theta}_{k}A^{\theta}_{k,k-1}\geq 0,\sum_{\theta}q^{\theta}_{k-1}A^{\theta}_{k,k-1}\leq 0,\sum_{\theta}q^{\theta}_{k-1}A^{\theta}_{k-1,k-2}\geq 0$ and, $\sum_{\theta}q^{\theta}_{k-2}A^{\theta}_{k-1,k-2}\leq 0$. Now consider,
@@
Thus, in 7, replacing the coefficients $\left(q^{\theta}_{k-1}\frac{q^{k-2}_{k-2}}{q^{k-2}_{k-1}}\right)$ with $q^{\theta}_{k-2}$ preserves the inequality, since this weakly decreases the coefficients of positive terms and weakly increases the coefficients of negative terms. Therefore,
@@
Then, in 10, replacing the coefficients $\left(q^{\theta}_{k-1}\frac{q^{k-1}_{k}}{q^{k-1}_{k-1}}\right)$ with $q^{\theta}_{k}$ preserves the inequality, since this weakly increases the coefficients of positive terms and weakly decreases the coefficients of negative terms. Therefore,
@@
Now suppose that $\{A^{\theta}_{k,l}\}$, where l<k2l<k-2, satisfies optimality. To complete the proof, it suffices to show that $\{A^{\theta}_{k,l-1}\}=\{A^{\theta}_{k,l}\}+\{A^{\theta}_{l,l-1}\}$ satisfies optimality. Again, $\{A^{\theta}_{l,l-1}\}$ satisfies optimality by construction. $\{A^{\theta}_{k,l}\}$ satisfies single crossing, so let $\theta_{n_{k,l}}$ be smallest $\theta$ such that $A^{\theta^{\prime}}_{k,l}\geq 0$ for every $\theta\geq\theta^{\prime}$. That is, $\theta_{n_{k,l}}$ is the crossing point for $A^{\theta^{\prime}}_{k,l}$. Consider,
@@
Finally, consider the system,
@@
There are exactly (M1)N(M-1)N independent columns in the coefficient matrix on the left hand side. [^19] Each row on the first matrix on the LHS has exactly two 1’s $N^{2}$ 0’s, where the positions of the 1’s ensure that the solution to the system induces the corresponding differences $A^{\theta}_{k,l}$. This is also the dimension of the vector on the RHS. So, the rank of the augmented matrix is same as that of the coefficient matrix and thus there exists a utility function $u(a,\theta)$ which induces the differences $\{A^{\theta}_{k,l}\}$. This completes the construction of a single crossing utility function that rationalizes the data, thereby completing the proof.
@@
5.2 Constructing the information structure and the choice rule